OpenStudy (anonymous):

Please help solving the differential equation: \[y'' = (y')^{3}\] Thanks

OpenStudy (anonymous):

\[y'' = (y')^{3}\]

OpenStudy (bahrom7893):

I think this is bernoulli..i think let me check..

OpenStudy (bahrom7893):

oh wow this looks so simple and I am not sure how to do this.. Well one solution is definitely: y=0, but that's trivial

OpenStudy (anonymous):

Haha, had the same first impression, I already thought about a substitution z=y', but then the particular Integral becomes a mess

OpenStudy (bahrom7893):

can u name the topic or something where I can look around.. I took ODEs last semestor..

OpenStudy (anonymous):

Unfortunately I cannot, have that one from a collection of miscellaneous exam problems.

OpenStudy (bahrom7893):

Ohh this is one of those substitutions.. when x is missing hold on.. and I meant *semester previously

OpenStudy (bahrom7893):

let me just find the section in the book..

OpenStudy (bahrom7893):

Reduction of order that's what it's called..

OpenStudy (bahrom7893):

okay there are two substitutions here..

OpenStudy (bahrom7893):

first one is: y' = z and y" = z' 2nd one is bernoulli..

OpenStudy (bahrom7893):

trying to solve this on paper..

OpenStudy (anonymous):

I tried it that way, when you then try to substitute z back to y you get an integral of the form \[\int\limits_{}^{}\sqrt{-1/(2x+C)}\] Is there no more elegant way?

OpenStudy (bahrom7893):

thinking.. trying this on paper..

OpenStudy (bahrom7893):

arghh im getting mixed up with all these variables.. y" = d^2y/dx^2 z' = dz/dx v' = dv/dx am I right?

OpenStudy (bahrom7893):

I have v' = -2 so far so what do I integrate it with respect to??

OpenStudy (anonymous):

Ok, you integrate this with respect to x, I finished of and got the result: y=\[C _{2} - \sqrt{2}-\sqrt{-x-C _{1}}\]

OpenStudy (bahrom7893):

actually i figured something out..

OpenStudy (bahrom7893):

let me just work it out and see what i have.

OpenStudy (anonymous):

I checked with wolfram alpha and that is correct, it's just, if I got this answer in an exam, I would be 100% sure to have made a mistake somewhere!

OpenStudy (anonymous):

Anyway bahrom, thanks for the effort!

OpenStudy (bahrom7893):

actually im almost done solving..

OpenStudy (bahrom7893):

okay done..

OpenStudy (bahrom7893):

let me take a picture and upload..

OpenStudy (bahrom7893):

here